OpenStudy (anonymous):

Solve for y. y – (–19) = 25 A. –44 B. –6 C. 6 D. 44

OpenStudy (anonymous):

@satellite73 @Ephemera

Vocaloid (vocaloid):

hint: y-(-19) = y + 19 y + 19 = 25 y = ?

OpenStudy (anonymous):

how would I figure this out sorry im tired

OpenStudy (anonymous):

im doing a timed test

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

ohhhhhhh I see it now its C. right?

OpenStudy (anonymous):

@ganeshie8

Vocaloid (vocaloid):

yes

OpenStudy (anonymous):

thanks